DEHYDROGENATION AND C-H BOND INSERTION OF PROPENE: La(η^{2}-C_{3}H_{4}) AND HLa(η^{3}-C_{3}H_{5})

\underline{SUDESH KUMARI} and DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Dehydrogenation and C-H bond insertion are observed in the reaction of laser-ablated La atoms with propene (C_{3}H_{6}) in a pulsed molecular beam source. Several dehydrogenated and inserted products are identified by the time-of-flight mass spectrometry. La(C_{3}H_{4}) formed from H_{2} elimination and HLa(C_{3}H_{5}) formed by C-H bond insertion are characterized by pulsed-field-ionization electron and ion spectroscopy, in combination with density functional theory. Two isomers of La(C_{3}H_{4}) are identified from 1,2- and 1,3-dehydrogenation. The adiabatic ionization energies of 1,2- and 1,3-dehydrogenated isomers are measured to be 40506(5) and 40941(5) cm^{−1}, respectively. For the inserted product HLa(C_{3}H_{5}), La atom is bound to the allyl radical in a three-fold binding mode (η^{3}). It is observed that the ionization energy of the HLa(η^{3}-C_{3}H_{5}) insertion complex (41130(5) cm^{−1}) is close to that of the 1,3-dehydrogented La(η^{2}-C_{3}H_{4}) species.